OpenStudy (anonymous):

If 120% of a is equal to 80% of b, which of the following is equal to a + b? A.)1.5a B.)2a C.)2.5a D.)3a E.)5a

OpenStudy (kira_yamato):

120% a = 80% b => 1.2 a = 0.8 b => b = 1.2a/0.8 = 1.5a a + b = a + 1.5a = 2.5a (C)